Let h : X −→ Y be defined by

h(x) :=



f(x) if x ∈ F
g
−1
(x) if x ∈ X − F

Now we must prove that h is injective and bijective. Starting with injectivity, let x1, x2 ∈
X such that h(x1) = h(x2). Assume x1 ∈ F and x2 ∈ X −F. Then h(x1) = f(x1) ∈ f(F)
and h(x2) = g
−1
(x2) ∈ g
−1
(X − F) = Y − f(F) as follows fDone Elements of set theory vol 3.pdf A Theorem 5.7 (Cantors-Beinstein Theorem). If X is equipotent to a subset of Y and Y irom question 2 above. Since
h(x1) = h(x2), and h(x1) = f(x1) ∈ f(F) and h(x2) ∈ Y − f(F), this is a contradiction
(question 3 why?).
Answer #1

Tot - : P(x) P(x) S.t . $(A): = x-914-f1A)) where f: X Y and giy-x are injective frenchon- They have alseady, broved, there iAlso, do t X-E h(a) = 91 (22) and x € X-F = gx2) E g*(X-F) h (262) - g 7182) E g(x-F) but gEX-F) = Y- f(F) = h (32) € Y- f (